OpenStudy (anonymous):

A certain type of chocolate comes in a package that's the shape of a triangular prism. How much chocolate is contained in this package? eighty-four cubic centimeters one hundred thirty-eight cubic centimeters one hundred sixty-eight cubic centimeters five hundred four cubic centimeters

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

The formula for the volume of a triangular prism is \(V = \dfrac{1}{2}lwh\)

OpenStudy (igreen):

Plug in what we know: \(V = \dfrac{1}{2}(14)(3)(4)\) Multiply: \(V = (7)(3)(4)\) \(V = (21)(4)\) \(V = 84\)

OpenStudy (anonymous):

thanks